The first-order decomposition of cyclopropane has a rate constant of 6.7 × 10-4 s-1. If the initial concentration of cyclopropane is 1.33 M, what is the concentration of cyclopropane after 644 s?
Answer #1

We know that for first order reaction

A= A0*e^-kt

here A= final concentration = ?

A0= initial concentration =1.33 M

k= rate constant = 6.7*10^-4 s^-1

t= time = 644 seconds

A= 1.33*e^-6.7*10^-4*644

A= 1.33*0.6495

A= 0.864 M